# Ecologically rational meta-learned inference explains human category learning

### v1

@@ I.2 Simulations
We simulated ERMI, MI, and PFN on the Johansen & Palmeri (2002) study for different betas values, from zero to one in steps of 0.1, for a total of 544 runs. The models interacted with each of the nine training stimuli 32 times with the ordering of the stimuli shuffled between runs. Predictions for the transfer stimuli were derived by concatenating them – one at a time – at the end of 32 training blocks in every run. By doing so, we were able to derive the model’s prediction for each unseen stimulus around $77$ times. In Figure 4 and 9, we reported average choice probabilities for the models using the $\beta$-term that minimized the pair-wise Euclidean distance between the human and model’s choice probabilities.
@@
kr-vs-kp classification 0.8257 0.8514 0.7986 0.8664 0.8450
credit-g classification 0.6421 0.6357 0.6350 0.6036 0.6150
diabetes classification 0.6771 0.7079 0.6786 0.6886 0.6950
spambase classification 0.5407 0.7664 0.7536 0.7993 0.7757
tic-tac-toe classification 0.5536 0.5950 0.6071 0.5914 0.6071
electricity classification 0.5543 0.6007 0.7036 0.6871 0.6436
pc4 Software defect prediction 0.7136 0.7521 0.7886 0.7707 0.7714
pc3 Software defect prediction 0.6514 0.7264 0.7357 0.7279 0.7107
kc2 Software defect prediction 0.5893 0.7314 0.7257 0.7257 0.7257
kc1 Software defect prediction 0.6271 0.6707 0.6743 0.6679 0.6521
pc1 Software defect prediction 0.5336 0.5964 0.6514 0.6064 0.6493
wdbc classification 0.9121 0.9207 0.9014 0.9221 0.9093
phoneme classification 0.5793 0.7314 0.6979 0.6921 0.7200
qsar-biodeg classification 0.5779 0.7014 0.6850 0.6921 0.7064
ilpd classification 0.5493 0.6386 0.6229 0.6121 0.6286
ozone-level-8hr classification 0.6614 0.6907 0.6707 0.6471 0.6950
banknote-authentication classification 0.7721 0.9229 0.8457 0.9657 0.9379
blood-transfution-service-center 0.4714 0.5493 0.5879 0.5671 0.6186
phishing websites classification 0.7929 0.8071 0.8157 0.8157 0.8129
bank-marketing classification 0.5829 0.5614 0.7386 0.7350 0.7171
wilt classification 0.5171 0.5736 0.6393 0.6371 0.6507
numerai28.6 classification 0.4857 0.4779 0.5029 0.4779 0.4986
churn classification 0.6321 0.7271 0.6800 0.7186 0.7329
@@
Appendix K Data and materials
We will make all experimental data and code for this project available under the following link: [RETRACTED].

### v2

@@ I.2 Simulations
We simulated ERMI, MI, and PFN on the Johansen & Palmeri (2002) study for different betas values, from zero to one in steps of 0.1, for a total of 544 runs. The models interacted with each of the nine training stimuli 32 times with the ordering of the stimuli shuffled between runs. Predictions for the transfer stimuli were derived by concatenating them – one at a time – at the end of 32 training blocks in every run. By doing so, we were able to derive the model’s prediction for each unseen stimulus around $77$ times. In Figure 4 and 13, we reported average choice probabilities for the models using the $\beta$-term that minimized the pair-wise Euclidean distance between the human and model’s choice probabilities.
@@
Abbreviations: Software defect prediction (SDP) and Service-Center (SC)
@@
kr-vs-kp 0.8257 0.8514 0.7986 0.8664 0.8450
credit-g 0.6421 0.6357 0.6350 0.6036 0.6150
diabetes 0.6771 0.7079 0.6786 0.6886 0.6950
spambase 0.5407 0.7664 0.7536 0.7993 0.7757
tic-tac-toe 0.5536 0.5950 0.6071 0.5914 0.6071
electricity 0.5543 0.6007 0.7036 0.6871 0.6436
pc4-sdp 0.7136 0.7521 0.7886 0.7707 0.7714
pc3-sdp 0.6514 0.7264 0.7357 0.7279 0.7107
kc2-sdp 0.5893 0.7314 0.7257 0.7257 0.7257
kc1-sdp 0.6271 0.6707 0.6743 0.6679 0.6521
pc1-sdp 0.5336 0.5964 0.6514 0.6064 0.6493
wdbc 0.9121 0.9207 0.9014 0.9221 0.9093
phoneme 0.5793 0.7314 0.6979 0.6921 0.7200
qsar-biodeg 0.5779 0.7014 0.6850 0.6921 0.7064
ilpd 0.5493 0.6386 0.6229 0.6121 0.6286
ozone-level-8hr 0.6614 0.6907 0.6707 0.6471 0.6950
banknote-authentication 0.7721 0.9229 0.8457 0.9657 0.9379
blood-transfution-sc 0.4714 0.5493 0.5879 0.5671 0.6186
phishing websites 0.7929 0.8071 0.8157 0.8157 0.8129
bank-marketing 0.5829 0.5614 0.7386 0.7350 0.7171
wilt 0.5171 0.5736 0.6393 0.6371 0.6507
numerai28.6 0.4857 0.4779 0.5029 0.4779 0.4986
churn 0.6321 0.7271 0.6800 0.7186 0.7329
@@
Appendix K Software and Data
We have made the data and code available under the following link: https://github.com/akjagadish/ermi